if Grace invested $6000 semi-annually in corporate bonds, beginning March 31 1990. If the yield is 8% compounded semi-annually, determine the amount of the bond on October 1st 1992
Answer by rfer(16322)

FV=P(1+r/n)^(n*t)
FV=6000(1+.08/2)^(2*2.5)
Fv=6000(1.04)^5
FV=6000(1.2166529)
FV=$7299.92